Directions : If y varies directly with x and y=7 when x=56,then find
Question: List the x-intercept and the y intercept of 2x+5y=10

Found 2 solutions by Boreal, ikleyn:
Answer by Boreal(15235)

y=kx
7=k*56
k=1/8
when x=0 y=2
(0, 2) is the y-intercept
when y=0 x=5 (5, 0) is the x-intercept.

Answer by ikleyn(52835)

.

In your post,  you  MISTAKENLY  mixed two different statements from two different problems.

Due to this reason,  your post   MAKES  NO  SENSE.

As a consequence,  the answer by @Boreal  makes no sense,  too.


Please,  be more attentive next time.


And NEVER combine more than one question/problem in your post (!) (!) (!)